\begin{thebibliography}{AHHM18}

%\bibitem[AH12]{AH12} {\sc H. Ando, U. Haagerup}, {\it Ultraproducts of von Neumann algebras.} J. Funct. Anal. {\bf 266} (2014), 6842--6913

%\bibitem[AHHM18]{AHHM18} {\sc H. Ando, U. Haagerup, C. Houdayer, A. Marrakchi}, {\it Structure of bicentralizer algebras and inclusions of type $\III$ factors.} Preprint (2018). {\tt arXiv:1804.05706}

%\bibitem[Ba93]{Ba93} {\sc L. Barnett}, {\it Free product von Neumann algebras of type ${\rm III}$.} Proc. Amer. Math. Soc. {\bf 123} (1995), 543--553.

%\bibitem[BIP18]{BIP18} {\sc R. Boutonnet, A. Ioana, J. Peterson}, {\it Properly proximal groups and their von Neumann algebras.} {\tt arXiv:1809.01881}

%\bibitem[BM86]{BM86} {\sc G. Baumslag, C.F. Miller III}, {\it Some Odd Finitely Presented Groups.} Bull. Lon. Math. Soc. {\bf 20} (1986), 239--244.

%\bibitem[BMO19]{BMO19} {\sc J. Bannon, A. Marrakchi, N. Ozawa}, {\it Full factors and co-amenable inclusions.} preprint.

%\bibitem[BO08]{BO08} {\sc N.P. Brown, N. Ozawa}, {\it C$^*$-algebras and finite-dimensional approximations.} Graduate Studies in Mathematics, {\bf 88}. American Mathematical Society, Providence, RI, 2008. xvi+509 pp.


%\bibitem[Bi89]{Bi89} {\sc D. Bisch}, {\it On the existence of central sequences in subfactors.} Trans. Amer. Math. Soc. {\bf 321} (1989), 117--128.

%\bibitem[BH16]{BH16} {\sc R. Boutonnet, C. Houdayer}, {\it Amenable absorption in amalgamated free product von Neumann algebras.} {To appear in Kyoto J. Math.} {\tt arXiv:1606.00808}

%\bibitem[Ch81]{Ch81} {\sc M. Choda}, {\it Inner amenability and fullness.} Proc. Amer. Math. Soc. {\bf 86} (1982), 663--666.

\bibitem[CE77]{CE77} {\sc M.-D. Choi and E. G. Effros}, {\it Injectivity and
operator spaces.} J.\ Funct.\ Anal.\  {\bf 24} (1977), 156--209.

%\bibitem[Co72]{Co72} {\sc A. Connes}, {\it Une classification des facteurs de type ${\rm III}$.} Ann. Sci. \'{E}cole Norm. Sup. {\bf 6} (1973), 133--252.

%\bibitem[Co74]{Co74} {\sc A. Connes}, {\it Almost periodic states and factors of type ${\rm III_1}$.} J. Funct. Anal. {\bf 16} (1974), 415--445.

%\bibitem[Co75a]{Co75a} {\sc A. Connes}, {\it Outer conjugacy classes of automorphisms of factors.} Ann. Sci. \'{E}cole Norm. Sup. {\bf 8} (1975), 383--419.

\bibitem[Co75]{Co75} {\sc A. Connes}, {\it Classification of injective factors. Cases ${\rm II_1}$, ${\rm II_\infty}$, ${\rm III_\lambda}$, $\lambda \neq 1$.} Ann. of Math. {\bf 74} (1976), 73--115.

%\bibitem[Co80]{Co80} {\sc A. Connes}, {\it A factor of type $\II_1$ with countable fundamental group.} J. Operator Theory {\bf 4} (1980), 151--153.

%\bibitem[Cor08]{Cor08} {\sc Y. Cornulier}, {\it Dense subgroups with Property (T) in Lie groups.} Comment. Math. Helv. {\bf 83}, 55--65, 2008.

%\bibitem[Co85]{Co85} {\sc A. Connes}, {\it Factors of type ${\rm III_1}$, property $L'_\lambda$ and closure of inner automorphisms.} J. Operator Theory {\bf 14} (1985), 189--211.

%\bibitem[CPS11]{CPS11} {\sc I. Chifan, S. Popa, J. O. Sizemore}, {\it Some OE- and W$^*$-rigidity results for actions by wreath product groups.} J. Funct. Anal. {\bf 263} (2012), no. 11, 3422--3448.

%\bibitem[CT76]{CT76} {\sc A. Connes, M. Takesaki}, {\it The flow of weights of factors of type ${\rm III}$.} Tohoku Math. J. {\bf 29} (1977), 473--575. 

%\bibitem[DTW19]{DTW19} {\sc B. Duchesne , R. Tucker-Drob, P. Wesolek}, {\it CAT(0) cube complexes and inner amenability.} {\tt arXiv:1903.01596}.

%\bibitem[Ef64a]{Ef64a} {\sc E. Effros}, {\it Convergence of closed subsets in a topological space}. Proc. Amer. Math. Soc. {\bf 16} (1965), no. 5, 929--931.

%\bibitem[Ef64b]{Ef64b} {\sc E. Effros}, {\it The Borel space of von Neumann algebras on a separable Hilbert space}. Paciﬁc. J. Math. {\bf 15} (1965), no. 4, 1153--1164.

%\bibitem[Ef73]{Ef73} {\sc E. G. Effros}, {\it Property Γ and inner amenability.} Proc. Amer. Math. Soc. {\bf 47} (1975), 483--486.


\bibitem[FK89]{FK89} {\sc H. Furstenberg, Y. Katznelson}, {\it Idempotents in compact semigroups and Ramsey theory.} Y. Israel J. Math. {\bf 68} (1989) 257--270




%\bibitem[GP03]{GP03} {\sc D. Gaboriau, S. Popa}, {\it An Uncountable Family of Non Orbit Equivalent Actions of $\F_n$.} J. Amer. Math. Soc. {\bf 18} (2005), 547--559.

%\bibitem[Ha73]{Ha73} {\sc U. Haagerup}, {\it The standard form of von Neumann algebras.} Math. Scand. {\bf 37} (1975), 271--283.

\bibitem[Ha85]{Ha85} {\sc U. Haagerup}, {\it Connes' bicentralizer problem and uniqueness of the injective factor of type ${\rm III_1}$.} Acta Math. {\bf 69} (1986), 95--148.

%\bibitem[Ho15]{Ho15} {\sc D.J. Hoff}, {\it Von Neumann algebras of equivalence relations with nontrivial one-cohomology.} J. Funct. Anal. {\bf 270} (2016), 1501--1536.

%\bibitem[Ho06]{Ho06} {\sc C. Houdayer}, {\it A new construction of factors of type ${\rm III_1}$.} J. Funct. Anal. {\bf 242} (2007), 375--399.

%\bibitem[Ho14]{Ho14} {\sc C. Houdayer}, {\it Gamma stability in free product von Neumann algebras.}  Commun. Math. Phys. {\bf 336} (2015), 831--851.

%\bibitem[HI15]{HI15} {\sc C. Houdayer, Y. Isono}, {\it Unique prime factorization and bicentralizer problem for a class of type ${\rm III}$ factors.} Adv. Math. {\bf 305} (2017), 402--455.

%\bibitem[HI15b]{HI15b} {\sc C. Houdayer, Y. Isono}, {\it Bi-exact groups, strongly ergodic actions and group measure space type ${\rm III}$ factors with no central sequence.} Comm. Math. Phys. {\bf 348} (2016), 991--1015.

%\bibitem[HI17]{HI17} {\sc C. Houdayer, Y. Isono}, {\it Factoriality, Connes' type III invariants and fullness of amalgamated free product von Neumann algebras} To appear in Proc. Roy. Soc. Edinburgh Sect. A. {\tt 	arXiv:1712.01747}.

%\bibitem[HMV16]{HMV16} {\sc C. Houdayer, A. Marrakchi, P. Verraedt}, {\it Fullness and Connes' $\tau$ invariant of type ${\rm III}$ tensor product factors.} To appear in J. Math. Pures Appl. {\tt arXiv:1611.07914}.

\bibitem[HP17]{HP17} {\sc C. Houdayer, S. Popa}, {\it Singular masas in type ${\rm III}$ factors and Connes' bicentralizer property.} Adv. Stud. in Pure Math. {\bf 80} (2019), 109--122.

%\bibitem[HR14]{HR14} {\sc C. Houdayer, S. Raum}, {\it Asymptotic structure of free Araki--Woods factors.} Math. Ann. {\bf 363} (2015), 237--267.

%\bibitem[HT18]{HT18} {\sc C. Houdayer, B. Trom}, {\it Structure of extensions of free Araki-Woods factors.} Preprint. {\tt 	arXiv:1812.08478}.

%\bibitem[HU15a]{HU15a} {\sc C. Houdayer, Y. Ueda}, {\it Asymptotic structure of free product von Neumann algebras.} \\ {\tt arXiv:1503:02460}.

%\bibitem[HU15]{HU15} {\sc C. Houdayer, Y. Ueda}, {\it Rigidity of free product von Neumann algebras.} Compos. Math. {\bf 152} (2016), 2461--2492.

%\bibitem[Io06]{Io06} {\sc A. Ioana}, {\it Rigidity results for wreath product $\rm II_1$ factors.} J. Funct. Anal. {\bf 252} (2007), no. 2, 763--791.

%\bibitem[Io08]{Io08} {\sc A. Ioana}, {\it Cocycle Superrigidity for Profinite Actions of Property (T) Groups.} Duke Math. J. {\bf 157} (2011), no. 2, 337--367

%\bibitem[Is16]{Is16} {\sc Y. Isono}, {\it On fundamental groups of tensor product $\II_1$ factors.} To appear in J. Inst. Math. Jussieu. {\tt arXiv:1608.06426.} 

%\bibitem[Is16b]{Is16b} {\sc Y. Isono}, {\it Cartan subalgebras of tensor products of free quantum group factors with arbitrary factors.} Anal. PDE {\bf 12} (2019), no. 5, 1295--1324.

%\bibitem[Is17]{Is17} {\sc Y. Isono}, {\it Unique prime factorization for infinite tensor product factors.} To appear in J. Funct. Anal. {\tt arXiv:1712.00925}.

%\bibitem[Is19]{Is19} {\sc Y. Isono}, {\it Unitary conjugacy for type $\rm III$ subfactors and W$^*$-superrigidity.} Preprint 2019. ArXiv:1902.01049.

%\bibitem[Jo81]{Jo81} {\sc V.F.R. Jones}, {\it Central sequences in crossed products of full factors.} Duke Math. J. {\bf 49} (1982), 29--33.

%\bibitem[KR97]{KR97} {\sc R.V. Kadison, J.R. Ringrose}, {\it Fundamentals of the theory of operator algebras. ${\rm I}$. Elementary theory.} Reprint of the 1983 original. Graduate Studies in Mathematics, {\bf 15}. American Mathematical Society, Providence, RI, 1997. xvi+398 pp.

%\bibitem[KST92]{KST92} {\sc Y. Kawahigashi, C.E. Sutherland, M. Takesaki}, {\it The structure of the automorphism group of an injective factor and the cocycle conjugacy of discrete abelian group actions.} Acta Math. {\bf 169} (1992), 105--130. 


%\bibitem[McD69]{McD69} {\sc D. McDuff}, {\it Central sequences and the hyperfinite factor.} Proc. London Math. Soc. {\bf 21} (1970), 443--461.

%\bibitem[Me82]{Me82} {\sc D. Meier}, {\it Non-hopfian groups.} J. London Math. Soc.(2) {\bf 26} (1982), 265--270.

%\bibitem[Ma73]{Ma73} {\sc O. Maréchal}, {\it Topologie et structure bor$\it \acute{e}$lienne sur l'ensembl des alg$\it \acute{e}$bres de von Neumann.} C. R. Acad. Sci. Paris S$\rm \acute{e}$r. A-B {\bf 276} (1973), 847--850.

%\bibitem[Ma16a]{Ma16a} {\sc A. Marrakchi}, {\it Solidity of type $\mathrm{III}$ Bernoulli crossed products.} To appear in Comm. Math. Phys. {\tt arXiv:1601.03666}

%\bibitem[Ma16]{Ma16} {\sc A. Marrakchi}, {\it Spectral gap characterization of full type $\mathrm{III}$ factors.} {To appear in J. Reine Angew. Math.} {\tt arXiv:1605.09613}

%\bibitem[Ma18]{Ma18} {\sc A. Marrakchi}, {\it Fullness of crossed products of factors by discrete groups.} Preprint. {\tt arXiv:1811.08274}.

\bibitem[Ma18]{Ma18} {\sc A. Marrakchi}, {\it Full factors, bicentralizer flow and approximately inner automorphisms.} Preprint. {\tt arXiv:1811.10253}.

%\bibitem[MvN43]{MvN43} {\sc F. Murray, J. von Neumann}, {\it Rings of operators.} ${\rm IV}$.  Ann. of Math. {\bf 44} (1943), 716--808.

%\bibitem[Oc85]{Oc85} {\sc A. Ocneanu}, {\it Actions of discrete amenable groups on von Neumann algebras.} Lecture Notes in Mathematics, {\bf 1138}. Springer-Verlag, Berlin, 1985. iv+115 pp.


%\bibitem[Oz02]{Oz02} {\sc N. Ozawa}, {\it There is no separable universal $\II_1$ factor.} Proc. Amer. Math. Soc. {\bf 132} (2004), 487--490.

%\bibitem[Oz03]{Oz03} {\sc N. Ozawa}, {\it Solid von Neumann algebras.} Acta Math.\ {\bf 192} (2004), 111--117.

%\bibitem[Oz16]{Oz16} {\sc N. Ozawa}, {\it A remark on fullness of some group measure space von Neumann algebras.} Compos. Math. {\bf 152} (2016), 2493--2502. 

\bibitem[Po81]{Po81} {\sc S. Popa}, {\it On a problem of R.V.\ Kadison on maximal abelian $\ast$-subalgebras in factors.} Invent. Math. {\bf 65} (1981), 269--281.



%\bibitem[Po83]{Po83} {\sc S. Popa}, {\it Hyperfinite subalgebras normalized by a given automorphism and related problems}, Proceedings of the Conference in Op. Alg. and Erg. Theory (Busteni 1983), Lect. Notes in Math., Springer-Verlag, {\bf 1132}, (1984) 421--433.

%\bibitem[Po86]{Po86}{\sc S. Popa}, {\it Correspondences}. INCREST preprint (1986).

%\bibitem[Po93]{Po93} {\sc S. Popa}, {\it Classification of Subfactors and Their Endomorphisms.} CBMS Regional Conference Series in Mathematics, {\bf 86}, (1995) 110 pp.

\bibitem[Po99]{Po99} {\sc On the relative Dixmier property for inclusions of $C^*$-algebras.} J. of Funct. Anal. {\bf 171} (2000), 139--154. 

\bibitem[Po01]{Po01} {\sc S. Popa}, {\it On a class of type ${\rm II_1}$ factors with Betti numbers invariants.} Ann. of Math. {\bf 163} (2006), 809--899.

\bibitem[Po03]{Po03} {\sc S. Popa}, {\it Strong rigidity of ${\rm II_1}$ factors arising from malleable actions of w-rigid groups $\rm I$.} Invent. Math. {\bf 165} (2006), 369--408.

%\bibitem[Po06]{Po06} {\sc S. Popa}, {\it On Ozawa's property for free group factors.} Int. Math. Res. Not. IMRN {\bf 2007}, no. 11, Art. ID rnm036, 10 pp.

%\bibitem[Po10]{Po10} {\sc S. Popa}, {\it On spectral gap rigidity and Connes' invariant $\chi(M)$.} Proc. Amer. Math. Soc. {\bf 138} (2010), 3531--3539.

%\bibitem[Po18]{Po18} {\sc S. Popa}, {\it On the vanishing cohomology problem for cocycle actions of groups on $\II_1$ factors.} Preprint. {\tt arXiv:1802.09964}.

%\bibitem[PV10]{PV10} {\sc S. Popa, S. Vaes}, {\it Actions of $\F_\infty$ whose $\II_1$ factors and orbit equivalence relations have prescribed fundamental group.} J. Amer. Math. Society {\bf 23} (2010), 383--403.


%\bibitem[PV11]{PV11} {\sc S. Popa, S. Vaes}, {\it On the fundamental group of $\II_1$ factors and equivalence relations arising from group actions.} Quanta of Maths, Proc. Conf. in honor of A.Connes' 60th birthday, Clay Math. Inst. Proc. 11 (2011), 519--541.

%\bibitem[Ra99]{Ra99} {\sc Y. Raynaud}, {\it On ultrapowers of non commutative $\rL_p$-spaces.} J. Operator Theory {\bf 48} (2002), 41--68.

\bibitem[Sc63]{Sc63} {\sc J. Schwartz}, {\it Two finite, non-hyperfinite, non-isomorphic factors.} Comm. Pure Appl. Math. {\bf 16} (1963), 19--26.

%\bibitem[Sh96]{Sh96} {\sc D. Shlyakhtenko}, {\it Free quasi-free states.} Pacific J. Math. {\bf 177} (1997), 329--368. 

\bibitem[Si15]{Si15} {\sc T. Sinclair}, {\it A very short proof of the existence of the injective enveloppe of an operator space.} \url{http://www.math.purdue.edu/~tsincla/injective-note-1.pdf}.

%\bibitem[Su80]{Su80} {\sc C. E. Sutherland}, {\it Cohomology and Extensions of von Neumann Algebras {\rm I}.} Publ. RIMS, Kyoto Univ.{\bf16} (1980), 105--133.




%\bibitem[Ta71]{Ta71} {\sc M. Takesaki}, {\it Conditional expectations in von Neumann algebras.} J. Funct. Anal. {\bf 9} (1972), 306--321.

%\bibitem[Ta03]{Ta03} {\sc M. Takesaki}, {\it Theory of operator algebras. ${\rm II}$.} Encyclopaedia of Mathematical Sciences, {\bf 125}. Operator Algebras and Non-commutative Geometry, 6. Springer-Verlag, Berlin, 2003. xxii+518 pp.

%\bibitem[Ta03b]{Ta03b} {\sc M. Takesaki}, {\it Theory of operator algebras. ${\rm III}$.} Encyclopaedia of Mathematical Sciences, {\bf 127}. Operator Algebras and Non-commutative Geometry, 8. Springer-Verlag, Berlin, 2003. xxii+548 pp.


%\bibitem[TU14]{TU14} {\sc R. Tomatsu, Y. Ueda}, {\it A characterization of fullness of continuous cores of type ${\rm III_1}$ free product factors.} Kyoto J. Math. {\bf 56} (2016), 599--610. 

%\bibitem[Ue10]{Ue10} {\sc Y. Ueda}, {\it Factoriality, type classification and fullness for free product von Neumann algebras.} Adv. Math. {\bf 228} (2011), 2647--2671.

%\bibitem[Va04]{Va04} {\sc S. Vaes}, {\it \'Etats quasi-libres libres et facteurs de type ${\rm III}$ (d'apr\`es D.\ Shlyakhtenko).} S\'eminaire Bourbaki, expos\'e 937, Ast\'erisque {\bf 299} (2005), 329--350.

%\bibitem[VV14]{VV14} {\sc S. Vaes, P. Verraedt}, {\it Classification of type ${\rm III}$ Bernoulli crossed products.} Adv. Math. {\bf 281} (2015), 296--332.

\end{thebibliography}

\documentclass[a4paper,11pt]{amsart}

\usepackage{amsmath, amsthm, amsfonts, amssymb}
\usepackage[bookmarks=false]{hyperref}
\usepackage{enumerate}

\usepackage{color}

\numberwithin{equation}{section}

\newtheorem*{mainthm}{Main theorem}

\newtheorem{thm}{Theorem}[section]
\newtheorem{lem}[thm]{Lemma}

\newtheorem{cor}[thm]{Corollary}
\newtheorem{prop}[thm]{Proposition}

\theoremstyle{definition}
\newtheorem{rem}[thm]{Remark}

\newtheorem{example}[thm]{Example}

\newtheorem{df}[thm]{Definition}

\newcommand{\C}{\mathbf{C}}
\newcommand{\Z}{\mathbf{Z}}

\newcommand{\B}{\mathbf{B}}

\newcommand{\Ad}{\operatorname{Ad}}

\newcommand{\III}{{\rm III}}

\begin{document}

\title[On the weak relative Dixmier property]{On the weak relative Dixmier property}

\address{Research Institute for Mathematical Sciences, Kyoto University, Kyoto, 606-8502 Japan}

\author{Amine Marrakchi}
\email{amine.marrakchi@math.u-psud.fr}

\thanks{The author is supported by JSPS}

\subjclass[2010]{46L10, 46L36, 46L40, 46L55}


\begin{abstract}
We show that every inclusion of von Neumann algebras with a faithful normal conditional expectation has the weak relative Dixmier property. This answers a question of Popa \cite{Po99}. The proof uses an improvement of Ellis' lemma for compact convex semigroup.
\end{abstract}

\maketitle






\section{Introduction}

Let $N \subset M$ be an inclusion of von Neumann algebras. Following \cite[Definition 1.1]{Po99}, we say that the inclusion $N \subset M$ has the \emph{weak relative Dixmier property} if for every $x \in M$, the weak$^*$ closed convex hull of $\{ uxu^* \mid u \in \mathcal{U}(N) \}$ intersects $N' \cap M$. A classical result of Dixmier asserts that this property is always satisfied when $N=M$.  However, as explained in \cite{HP17}, if one does not make any assumption on the inclusion $N \subset M$, then the weak relative Dixmier property is not necessarily satisfied and can be very subtle as demonstrated by the following two examples:
 \begin{itemize}
 \item The inclusion $M \subset \B(H)$ has the weak relative Dixmier property if and only if $M$ is AFD. The if direction was first noticed by Schwartz \cite{Sc63} while the converse follows from Connes' celebrated work \cite{Co75}. 
 \item A type $\III_1$ factor $M$ has a trivial bicentralizer if and only if the inclusion $N \subset M$ has the weak relative Dixmier property, where $N$ is the \emph{continuous core} of $M$. This equivalence is due to Haagerup \cite{Ha85}. It is still a deep open problem to determine whether every type $\III_1$ factor has a trivial bicentralizer.
 \end{itemize}

On the other hand, it is an elementary fact \cite[Lemma 1.2]{Po99} that an inclusion $N \subset M$ has the weak relative Dixmier property if $M$ is tracial or more generally if there exists a faithful normal state $\varphi$ on $M$ such that $N \subset M_\varphi :=\{a \in M \mid \forall x \in M, \:  \varphi(a x)=\varphi( x a) \}$. This kind of averaging argument is actually a fundamental tool in von Neumann algebra theory. For example, it plays a key role in \cite{Po81} or in Popa's intertwining theory \cite{Po01, Po03}. In \cite[Remark 1.5]{Po99} (see also \cite[Problem 7]{HP17}), it was suggested that the weak relative Dixmier property might more generally hold for any inclusion \emph{with expectation} (this would in particular cover \cite[Corollary 1.4]{Po99}). Our main theorem solves this question.

\begin{mainthm} \label{main}
Let $N \subset M$ be an inclusion of von Neumann algebras. Suppose that there exists a faithful normal state $\varphi$ on $M$ such that $N$ is the range of a $\varphi$-preserving conditional expectation $E : M \rightarrow N$. Then the inclusion $N \subset M$ has the weak relative Dixmier property.
\end{mainthm}
The difficulty in the proof of this theorem arises only when $N$ is of type $\III_1$. Indeed, in all other cases, one can use discrete decompositions in order to reduce the problem to the tracial case. To deal with the remaining type $\III_1$ case, we use a recent result of the author \cite{Ma18} asserting the existence of an irreducible AFD subfactor with expectation inside $N$. We combine it with a new averaging argument for compact convex semigroups based on Ellis' lemma and inspired by Sinclair's proof of the existence of the injective enveloppe \cite{Si15}.

\section{Compact convex semigroups}
\begin{df}
A \emph{compact semigroup} $K$ is a non-empty compact space equipped with a semigroup operation such that $K \ni a \mapsto ab \in K$ is continuous for every given $b \in K$.
\end{df} 

The proof of the following result, known as \emph{Ellis' lemma}, is very short but it has many surprising applications to combinatorics \cite{FK89}. Recently, Sinclair applied it to give a very short proof of the existence of the injective enveloppe of an operator system \cite{Si15}.

Recall that an element $e$ of a semigroup $S$ is \emph{idempotent} if $ee=e$. We define an order relation between idempotents as follows: $e < f$ if and only if $ef=fe=e$.

\begin{thm}[\cite{FK89}] \label{Ellis} Let $K$ be a compact semigroup. Then $K$ contains an idempotent which is minimal for the order relation $<$.
\end{thm}
\begin{rem}
It is also shown in \cite{FK89} that an idempotent $e \in K$ is minimal if and only if $Ke$ is a minimal closed left ideal.
\end{rem}

\begin{df}
A \emph{compact convex semigroup} $K$ is a non-empty compact convex space (in some locally convex topological vector space) equipped with a semigroup operation such that $K \ni a \mapsto ab \in K$ is affine and continuous for every given $b \in K$.
\end{df}

\begin{example}
Let $M$ be a von Neumann algebra and let $\mathrm{UCP}(M)$ be the set of all UCP maps from $M$ into $M$, equipped with the topology of pointwise weak$^*$ convergence. Then $\mathrm{UCP}(M)$ is a compact convex semigroup.
\end{example}

We now strengthen the conclusion of Theorem \ref{Ellis} for compact convex semigroups.

\begin{thm} \label{minimal idempotent}
Let $K$ be a compact convex semigroup. Let $e \in K$ be a minimal idempotent. Then we have $exe=e$ for all $x \in K$.
\end{thm}
\begin{proof}
Let $I=Ke$ be the minimal closed left ideal generated by $e$. Take any $x \in K$ and let $J=\{ \frac{1}{2}(bxe+b) \mid b \in I \}$. Then $J \subset I$ and $J$ is a closed left ideal of $K$. By minimality of $I$, we must have $J=I$. Let $a$ be an extremal point of the convex compact space $I$. Since $J=I$, we can find $b \in I$, such that $a=\frac{1}{2}(bxe+b)$. Because $a$ is extremal, this forces $a=bxe=b$. Thus $axe=a$. Since this holds for every extremal point $a \in I$, and since $I$ is the closed convex hull of its extremal points (Krein-Millmann theorem), we conclude that $axe=a$ for all $a \in I$. In particular, we have $exe=e$ as we wanted.
\end{proof}

Recall that if $A$ is a unital $C^*$-algebra and $\Phi : A \rightarrow A$ is an idempotent UCP map, then $\Phi(A)$ equipped with the Choi-Effros product $(x,y) \mapsto \Phi(xy)$ is a $C^*$-algebra \cite[Theorem 3.1]{CE77}. By combining this fact with Theorem \ref{minimal idempotent}, one can associate to each compact convex semigroup of UCP maps on a von Neumann algebra a canonical ``boundary" $C^*$-algebra. See \cite{Si15} for examples.

\begin{cor}
Let $M$ be a von Neumann and $K \subset \mathrm{UCP}(M)$ a closed convex semigroup. Let $\Phi$ and $\Psi$ be two minimal idempotents. Then $\Psi|_{\Phi(M)} : \Phi(M) \rightarrow \Psi(M)$ is an isomorphism of $C^*$-algebras with inverse $\Phi|_{\Psi(M)}$.
\end{cor}
\begin{proof}
By Theorem \ref{minimal idempotent}, we have $\Psi \circ \Phi \circ \Psi=\Psi$ and $\Phi \circ \Psi \circ \Phi=\Phi$. Terefore $\Psi|_{\Phi(M)}$ and $\Phi|_{\Psi(M)}$ are UCP maps which are inverse to each others. Thus they must be isomorphisms of $C^*$-algebras.
\end{proof}

\begin{prop} \label{faithful idempotent}
Let $M$ be a von Neumann and $K \subset \mathrm{UCP}(M)$ a closed convex semigroup. Let $M^K=\{ x \in M \mid \forall \Psi \in K, \: \Psi(x)=x \}$. Let $\Phi \in K$ be a minimal idempotent. If $\Phi$ is faithful, then $M^K$ is a $C^*$-subalgebra of $M$ and $\Phi$ is a conditional expectation of $M$ onto $M^K$.
\end{prop}
\begin{proof}
For every $x \in \Phi(M)$, we have $\Phi(x^*x) \geq \Phi(x^*)\Phi(x)=x^*x$. But we have $\Phi(\Phi(x^*x) - x^*x) =0$. Thus $\Phi(x^*x)=x^*x$ because $\Phi$ is faithful. In particular, $x^*x \in \Phi(M)$ for all $x \in \Phi(M)$. By the polarization identity, this means that $\Phi(M)$ is a subalgebra of $M$. Then $\Phi$ is a faithful conditional expectation onto $\Phi(M)$. Now, take a unitary $u \in \Phi(M)$. For every $\Psi \in K$, we have $\Phi(\Psi(u))=u$ by Theorem \ref{minimal idempotent}. Since $\Phi$ is a faithful conditional expectation, this forces $\Psi(u)=u$. Thus $u \in M^K$. Since this holds for every unitary $u \in \Phi(M)$, we conclude that $\Phi(M)=M^K$.
\end{proof}

\section{The weak relative Dixmier property}

Let $N \subset M$ be an inclusion of von Neumann algebras. Let $\mathcal{D}(N,M) \subset \mathrm{UCP}(M)$ be the closed convex semigroup generated by $\Ad(u)$ for all $u \in \mathcal{U}(N)$. Note that $M^{\mathcal{D}(N,M)}=N' \cap M$. Observe that the inclusion $N \subset M$ has the weak relative Dixmier property if and only if $\mathcal{D}(N,M)$ contains a conditional expectation onto $N' \cap M$.

We need the following key lemma which relies on \cite[Theorem D]{Ma18}.
\begin{lem} \label{dixmier III1}
Let $N$ be a type $\III_1$ factor with separable predual. Let $\varphi$ be a faithful normal state on $N$. Then $\varphi \in \mathcal{D}(N,N)$.
\end{lem}
\begin{proof}
By \cite[Theorem D]{Ma18}, $N$ contains an irreducible subfactor with expectation $P \subset N$ such that $P$ is AFD of type $\III_1$. Let $\psi$ be a faithful normal state on $N$ such that $P$ is globally invariant by $\sigma^\psi$ and $P_\psi' \cap P=\C$. Since $P$ is AFD and irreducible in $N$, we know that $\mathcal{D}(P,N)$ contains a state $\phi$. Note that $\phi=\phi \circ E^\psi_P$. Thus, we have $\phi \circ E^\psi_{P_\psi' \cap N}=\psi$. Since $P_\psi$ centralizes $\psi$, we have $E^\psi_{P_\psi' \cap N} \in \mathcal{D}(P_\psi,N)$. We conclude that $\psi=\phi \circ E^\psi_{P_\psi' \cap N} \in \mathcal{D}(P,N) \subset \mathcal{D}(N,N)$. Finally, by Connes-St\o rmer transitivity theorem, we get $\varphi \in \mathcal{D}(N,N)$. 
\end{proof}

We are now ready to prove our main theorem.

\begin{proof}[Proof of the main theorem] First note that we can always write $N$ as an increasing union of a net of subalgebras $(N_i)_{i \in I}$ with separable predual which are globally invariant under $\sigma^\varphi$. Thus, we may assume that $N$ has a separable predual. Then, we may also assume that $M$ has a separable predual. Since $\mathcal{Z}(N)$ has the weak relative dixmier property in $M$, we may replace $M$ by $\mathcal{Z}(N)' \cap M$ and reduce to the case where $\mathcal{Z}(N) \subset \mathcal{Z}(M)$. Now by using the desintegration theory, we may assume that $N$ is a factor. If $N$ is semifinite, the result is already known. So we may assume that $N$ is a type $\III$ factor. If $N$ is not of type $\III_1$, then $N \cong N_0 \rtimes \Z$ for some semifinite von Neumann algebra with expectation $N_0 \subset N$. Since $N_0$ has the weak relative dixmier property in $M$ and $\Z$ is amenable, then $N$ also has the weak relative dixmier property in $M$. 

Finally, we deal with the case where $N$ is of type $\III_1$. By compactness of $\mathcal{D}(N,M)$, any element of $\mathcal{D}(N,N)$ can be extended to an element of $\mathcal{D}(N,M)$. In particular, thanks to Lemma \ref{dixmier III1}, we can find $\Psi \in \mathcal{D}(N,M)$ such that $\Psi |_N=\varphi$. Take $\Phi_0$ a minimal idempotent in $\mathcal{D}(N,M)$ and let $\Phi=\Phi_0 \circ \Psi$ which is again a minimal idempotent of $\mathcal{D}(N,M)$. Observe that $E$ commutes with every element of $\mathcal{D}(N,M)$. Thus, we have 
$$ E \circ \Phi = \Phi \circ E= \Phi_0 \circ \Psi \circ E=\Phi_0 \circ \varphi= \varphi.$$
In particular, $\Phi$ is faithful and by Proposition \ref{faithful idempotent}, we conclude that $\Phi$ is a conditional expectation on $N' \cap M$ (in fact $\Phi$ is the unique $\varphi$-preserving conditional expectation onto $N' \cap M$).
\end{proof}

\begin{rem}
Let $M \subset \B(H)$ be a von Neumann algebra. Let $\Phi \in \mathcal{D}(M,\B(H))$ be a minimal idempotent. Then $A_\Phi=\Phi(\B(H))$ (whose isomorphism class as a $C^*$-algebra does not depend on the choice of $\Phi$) contains $M'$. We have $A_\Phi=M'$ if and only if $M$ is injective. In general, $A_\Phi$ contains the injective enveloppe of $M'$ but it is not clear if they are equal or not. Notice that if $M \subset N \subset \B(H)$, then $M \subset N$ has the weak relative Dixmier property if and only if $A_\Phi \cap N=M' \cap N$.
\end{rem}

\bibliographystyle{plain}

\end{document}